# Shaocheng Liu and Qi Chen

### v2

@@ IV-D Entropy functions on the faces involving partitions
In this subsection, we characterize the faces where the distribution of the entropy function’s characterizing random vector involves number partitions. A $t$-partition of a positive $k$ with $1\leq t\leq k$ is a $t$-tuple $\bm{\alpha}=(\alpha_{1},\alpha_{2},\dots,\alpha_{t})$ with $\alpha_{i}$ positive integer and $\sum_{i=1}^{t}\alpha_{i}=k$. The family of all such partitions of $k$ is denoted by $\mathcal{P}(k)$ and $H(\bm{\alpha})$ reprensent the shannon entropy $H(\frac{\alpha_{i}}{k}:i\in N_{t})$.
@@
For $F=(\mathcal{W}_{2}^{12},U_{2,3}^{134})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log{k},a=H(\dfrac{\alpha_{1}}{k},\dfrac{\alpha_{2}}{k},\dots,\dfrac{\alpha_{t}}{k})$ , where integer $k>0$ and $(\alpha_{1},\alpha_{2},\dots,\alpha_{t})$ is a partition of $k$.
$a$$b$1111222233334444Olog2log3log4log2log3log41:(0.92,0.67):10.920.671:(0.92,0.67)1 : ( 0.92 , 0.67 )$2:(0.81,1.19)$$3:(1.00,1.00)$$4:(1.50,0.50)$
@@
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,3,4\},|K|=2,\{i,j\}=N_{4}\setminus K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
By (IV.D.17)-(IV.D.20), $X_{1},X_{3}$ and $X_{4}$ are pairwise independent and each one is a function of the other two. Thus, $(X_{1},X_{3},X_{4})$ is the characterizing random vector of an entropy function on $U_{2,3}$, an extreme ray of $\Gamma_{3}$, and $X_{i}$ is uniformly distributed on $\mathcal{X}_{i}$ and $H(X_{i})=\log k$ where $|\mathcal{X}_{i}|=k$ for $i=1,3,4$.
@@
$H\left(X_{3}\right)$ $=a+b=\log{k},$ (IV.D.24)
$H\left(X_{4}\right)$ $=a+b=\log{k}.$ (IV.D.25)
@@
By (IV.D.27), $X_{2}$ is a function of $X_{1}$. Thus there exists a unique $x_{2}\in\mathcal{X}_{2}$ with $p(x_{1},x_{2})>0$, then $\mid\mathcal{X}_{2}\mid\leq\mid\mathcal{X}_{1}\mid=k$ and
$p(x_{2})=\sum\limits_{x_{1}}p(x_{1},x_{2})=\dfrac{\alpha\left(x_{2}\right)}{k},$ (IV.D.28)
where $\alpha(x_{2})$ is the number of $x_{1}$ with $p(x_{1},x_{2})>0$. Then $(\alpha\left(x_{2}\right):x_{2}\in\mathcal{X}_{2})$ is a number partition of $k$, i.e.,$\sum\limits_{x_{2}\in\mathcal{X}_{2}}\alpha(x_{2})=k$. By (IV.D.23), $a$ can only take the vaule $H(\dfrac{\alpha(x_{2})}{k}:x_{2}\in\mathcal{X}_{2})$.
Now we prove all these points are entropic. Let $X_{1},X_{3}$ be indenpendent and uniformly distributed on $\mathcal{X}_{1}=\mathcal{X}_{3}=\{0,1,\dots,k-1\}$. Let $X_{4}=X_{1}+X_{3}$ mod $k$. It can be checked that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}=\{0,1,\dots,k-1\}$. Let $\left(A(x_{2}):x_{2}\in\mathcal{X}_{2}\right)$ be a partition of $\{0,1,\dots,k-1\}$, where $|A(x_{2})|=\alpha(x_{2}).$ Let $X_{2}$ be a function $\varphi$ of $X_{1}$, where $\varphi(x_{1})=x_{2}$ if $x_{1}\in A(x_{2})$. Then it can be checked that $(X_{i},i\in N_{4})$ is the characterizing random vector of $\mathbf{h}\in F$. ∎
@@ IV-E Entropy functions on the faces involving $U_{3,4}$
In this subsection, we characterize entropy functions on the $5$ $2$-dimensional faces of $\Gamma_{4}$ with one of its extreme rays $U_{3,4}$. As introduced in [6], uniform matroid $U_{3,4}$ corresponding to Latin cubes or orthogonal arrays $\mathrm{OA}(3,4)$, the characterizing random vectors of these entropy functions are distributed on the combinatorial structures extending Latin cubes.
@@
For $F=(U_{3,4},U^{1}_{1,1})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a$= $\log{k}$ for integer $k>0$
@@
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
Note that $X_{3}$ and $X_{4}$ are independent, by Lemma 1, $X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and so $H(X_{3})=H(X_{4})=\log k$ where $k=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$.
For the “if ” part, it can be proved by Lemma 2 and the fact that $a=\log{k}$ on the ray $U_{3,4}$, and whole ray $U^{1}_{1,1}$ are entropic. ∎
@@
For $F=(U_{3,4},U^{12}_{1,2})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a$= $\log{k}$ for integer $k>0$
@@
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),\{i,j\}\neq\{1,2\},i,j,k\in N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
By (IV.E.32), $X_{3}$ and $X_{4}$ are independent. Due to Lemma 1, we conclude that $X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and so $H(X_{3})=H(X_{4})=\log k$ where $k=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$.
On the other hand, the “if” part of the theorem can be proved by Lemma 2 and the fact that $a=\log{k}$ on the ray $U_{3,4}$ and whole ray $U^{12}_{1,2}$ are entropic. ∎
@@
For $F=(U_{3,4},U^{123}_{2,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b$= $\log{k}$ for integer $k>0$
$a$$b$Olog2log3log4log2log3log4
@@
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
By (IV.E.49), (IV.E.50) and (IV.E.52), $X_{1}$, $X_{2}$ and $X_{3}$ are pairwise independent. Then by Lemma 1, we conclude that $X_{i}$ are uniformly distributed on $\mathcal{X}_{i}$ for $i\in N_{3}$ and $H(X_{1})=H(X_{2})=H(X_{3})=\log k$ where $k=|\mathcal{X}_{1}|=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|$.
Now we prove the “if ” part. Let $X_{2}$, $X_{3}$ and $X_{4}$ be mutually independent and distributed on $\{0,1,2,...,k-1\}$, where $X_{2}$ and $X_{3}$ are uniform and $H(X_{4})=a$ for some $a\leq\log k$. Let $X_{1}=X_{2}+X_{3}+X_{4}\mod k$. Then for any $i\in\mathcal{X}_{1}$,
$p_{X_{1}}(i)=\sum_{x_{2},x_{3},x_{4}:\atop x_{2}+x_{3}+x_{4}\equiv i\mod k}p_{X_{2},X_{3},X_{4}}(x_{2}.x_{3},x_{4}).$ (IV.E.68)
@@
$p_{X_{1}}(i)$ $=\sum_{j=0}^{k-1}\sum_{m=0}^{k-1}\dfrac{1}{k}p_{X_{3}}(m)p_{X_{4}}(j)$ (IV.E.72)
$=\dfrac{1}{k}\sum_{j=0}^{k-1}p_{X_{4}}(j)\sum_{m=0}^{k-1}p_{X_{3}}(m)$ (IV.E.73)
$=\dfrac{1}{k},$ (IV.E.74)
which implies $X_{1}$ is uniformly distributed on $\mathcal{X}_{1}$ and $H(X_{1})=\log k$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,b)$. ∎
@@
For $F=(U_{3,4},U^{123}_{1,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a$= $\log{k}$ for integer $k>0$
@@
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),k\neq 4,\text{ distinct }i,j,k\in N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
It holds for any $x_{4}^{\prime}\in\mathcal{X}_{4}$, which implies that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}$ , and so $H(X_{4})=\log k$ where $k=|\mathcal{X}_{4}|$.
On the other hand, the “if ” part of the theorem can be proved by Lemma 2 and the fact that $a=\log{k}$ on the ray $U_{3,4}$ and whole ray $U^{123}_{1,3}$ are entropic. ∎
@@
For $F=(U_{3,4},U_{1,4})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a\geq 0,b>0$ or $(a,b)=(\log{k},0)$ for some positive integer $k$.
@@
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
@@
Consider the tripartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}$ and $(x_{i},x_{j})\in E$ if and only if $p(x_{i},x_{j})>0$, ${i,j}\in N_{3},i\neq j$. Assume $G$ has $t$ connected components. We denote the number of the vertices in $\mathcal{X}_{i}$ by $n_{i}^{(j)},i=1,2,3,j=1,2,\cdots,t$ in the connected component $C_{j}$ and the probability mass of $C_{j}$ by $p_{j}$. By Lemma 3, we obtain $n_{1}^{(j)}=n_{2}^{(j)}=n_{3}^{(j)}$ and so it can be simplified to $n^{(j)}$. The probability mass of the vertices, the edges and the triangles in $C_{j}$ are
@@
By Lemma 2 and the fact that $a=\log k$ on the ray $U_{3,4}$, and the whole ray $U_{1,4}$ are entropic, all $\mathbf{h}=(a,b)\in F$ are entropic when $a=\log k$ for positive integer $k$, and $b\geq 0$. In the following, we only need to show that for all $\mathbf{h}=(a,b)\in F$ are entropic, when $a>0,b>0$ with $a\neq\log k$ for some integer $k>0$.
Consider $a\neq\log{k},b>0$. Let $c=\lfloor 2^{a}\rfloor$. Let $n^{(j)}=c$ for all $j\in N_{t-1}$, $n^{(t)}=c+l$ for some sufficiently large $l$ and $10<p_{i}<$ 1, $i=1,2,\dots,k$ be positive values satisfying $H(p_{1},...,p_{t})=b$ such that
@@
$=(1-p_{t})\log c+p_{t}\log{(c+l)}=a.$ (IV.E.145)
Let $X^{(j)}_{1}$ ,$X^{(j)}_{2}$ and $X^{(j)}_{3}$ be mutually independent and uniformly distributed on $N^{\prime}_{c}\triangleq\{0,1,\dots,c-1\}$ , and $X^{(j)}_{4}=X^{(j)}_{1}+X^{(j)}_{2}+X^{(j)}_{3}\mod c$ for $j\in N_{t-1}$. Let $X^{(t)}_{1}$, $X^{(t)}_{2}$ and $X^{(t)}_{3}$ be mutually independent and uniformly distributed on $N^{\prime}_{c+l}$ and $X^{(t)}_{4}=X^{(t)}_{1}+X^{(t)}_{2}+X^{(t)}_{3}\mod(c+l)$. Assume $X^{(j)}_{i}$ is distributed on $\mathcal{X}^{(j)}_{i}$ for $i\in N_{4}$ and $j\in N_{t}$. Let $X_{i},i\in N_{4}$ be distributed on $N^{\prime}_{tc+l}$ such that
$p_{X_{N_{4}}}(x_{i}+(j-1)c:i\in N_{4})=p_{j}p_{X^{(j)}_{1},X^{(j)}_{2},X^{(j)}_{3},X^{(j)}_{4}}(x_{1},x_{2},x_{3},x_{4})$ (IV.E.146)
for $j\in N_{t}$ and $(x_{1},x_{2},x_{3},x_{4})\in\mathcal{X}^{(j)}_{N_{4}}$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,b)$. Hence, all $\mathbf{h}=(a,b)\in F$ are entropic when $a>0,b>0$.
Note that $a=\log k$ on the ray $U_{3,4}$ and the whole ray $U_{1,4}$ are entropic. The proof is accomplished. ∎

### v3

@@ IV-D Entropy functions on the faces involving partitions
In this subsection, we characterize the faces where the distribution of the entropy function’s characterizing random vector involves number partitions. A $t$-partition of a positive $v$ with $0\leq t\leq v-1$ is a $t$-tuple $\bm{\alpha}=(\alpha_{0},\alpha_{1},\dots,\alpha_{t-1})$ with $\alpha_{i}$ positive integer and $\sum_{i=0}^{t-1}\alpha_{i}=v$. The family of all such partitions of $v$ is denoted by $\mathcal{P}(v)$ and $H(\bm{\alpha})$ reprensent the shannon entropy $H(\frac{\alpha_{i}}{v}:i\in\mathbb{I}_{t})$.
@@
For $F=(\mathcal{W}_{2}^{12},U_{2,3}^{134})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log{v},a=H(\dfrac{\alpha_{1}}{v},\dfrac{\alpha_{2}}{v},\dots,\dfrac{\alpha_{t}}{v})$ , where integer $v>0$ and $(\alpha_{0},\alpha_{1},\dots,\alpha_{t-1})$ is a partition of $v$.
[…]
@@
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,3,4\},$
$\qquad\qquad\qquad|K|=2,\{i,j\}=N_{4}\setminus K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
By (IV.D.17)-(IV.D.20), $X_{1},X_{3}$ and $X_{4}$ are pairwise independent and each one is a function of the other two. Thus, $(X_{1},X_{3},X_{4})$ is the characterizing random vector of an entropy function on $U_{2,3}$, an extreme ray of $\Gamma_{3}$, and $X_{i}$ is uniformly distributed on $\mathcal{X}_{i}$ and $H(X_{i})=\log v$ where $|\mathcal{X}_{i}|=v$ for $i=1,3,4$.
@@
$H\left(X_{3}\right)$ $=a+b=\log{v},$ (IV.D.24)
$H\left(X_{4}\right)$ $=a+b=\log{v}.$ (IV.D.25)
@@
By (IV.D.27), $X_{2}$ is a function of $X_{1}$. Thus there exists a unique $x_{2}\in\mathcal{X}_{2}$ with $p(x_{1},x_{2})>0$, then $\mid\mathcal{X}_{2}\mid\leq\mid\mathcal{X}_{1}\mid=v$ and
$p(x_{2})=\sum\limits_{x_{1}}p(x_{1},x_{2})=\dfrac{\alpha\left(x_{2}\right)}{v},$ (IV.D.28)
[…]
Now we prove all these points are entropic. Let $X_{1},X_{3}$ be indenpendent and uniformly distributed on $\mathcal{X}_{1}=\mathcal{X}_{3}=\mathbb{I}_{v}$. Let $X_{4}=X_{1}+X_{3}$ mod $v$. It can be checked that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}=\mathbb{I}_{v}$. Let $\left(A(x_{2}):x_{2}\in\mathcal{X}_{2}\right)$ be a partition of $\mathbb{I}_{v}$, where $|A(x_{2})|=\alpha(x_{2}).$ Let $X_{2}$ be a function $\varphi$ of $X_{1}$, where $\varphi(x_{1})=x_{2}$ if $x_{1}\in A(x_{2})$. Then it can be checked that $(X_{i},i\in N_{4})$ is the characterizing random vector of $\mathbf{h}\in F$. ∎
@@ IV-E Entropy functions on the faces involving $U_{3,4}$
In this subsection, we characterize entropy functions on the 55 22-dimensional faces of $\Gamma_{4}$ with one of its extreme rays $U_{3,4}$. As introduced in [7], uniform matroid $U_{3,4}$ corresponding to Latin cubes or orthogonal arrays $\mathrm{OA}(3,4)$, the characterizing random vectors of these entropy functions are distributed on the combinatorial structures extending Latin cubes.
@@
For $F=(U_{3,4},U^{1}_{1,1})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
@@
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
Note that $X_{3}$ and $X_{4}$ are independent, by Lemma 1, $X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and so $H(X_{3})=H(X_{4})=\log v$ where $v=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$.
For the “if” part, it can be proved by Lemma 2 and the fact that $a=\log{v}$ on the ray $U_{3,4}$, and whole ray $U^{1}_{1,1}$ are entropic. ∎
@@
For $F=(U_{3,4},U^{12}_{1,2})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
@@
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),\{i,j\}\neq\{1,2\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
By (IV.E.32), $X_{3}$ and $X_{4}$ are independent. Due to Lemma 1, we conclude that $X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and so $H(X_{3})=H(X_{4})=\log v$ where $v=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$.
On the other hand, the “if” part of the theorem can be proved by Lemma 2 and the fact that $a=\log{v}$ on the ray $U_{3,4}$ and whole ray $U^{12}_{1,2}$ are entropic. ∎
@@
For $F=(U_{3,4},U^{123}_{2,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if a+ba+b= $\log{v}$ for integer $v>0$ .
@@
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
By (IV.E.49), (IV.E.50) and (IV.E.52), $X_{1}$, $X_{2}$ and $X_{3}$ are pairwise independent. Then by Lemma 1, we conclude that $X_{i}$ are uniformly distributed on $\mathcal{X}_{i}$ for $i\in N_{3}$ and $H(X_{1})=H(X_{2})=H(X_{3})=\log v$ where $v=|\mathcal{X}_{1}|=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|$.
Now we prove the “if ” part. Let $X_{2}$, $X_{3}$ and $X_{4}$ be mutually independent and distributed on $\mathbb{I}_{v}$, where $X_{2}$ and $X_{3}$ are uniform and $H(X_{4})=a$ for some $a\leq\log v$. Let $X_{1}=X_{2}+X_{3}+X_{4}\mod v$. Then for any $i\in\mathcal{X}_{1}$,
$p_{X_{1}}(i)=\sum_{x_{2},x_{3},x_{4}:\atop x_{2}+x_{3}+x_{4}\equiv i\mod v}p_{X_{2},X_{3},X_{4}}(x_{2}.x_{3},x_{4}).$ (IV.E.68)
@@
$p_{X_{1}}(i)$ $=\sum_{j=0}^{v-1}\sum_{m=0}^{v-1}\dfrac{1}{v}p_{X_{3}}(m)p_{X_{4}}(j)$ (IV.E.72)
$=\dfrac{1}{v}\sum_{j=0}^{v-1}p_{X_{4}}(j)\sum_{m=0}^{v-1}p_{X_{3}}(m)$ (IV.E.73)
$=\dfrac{1}{v},$ (IV.E.74)
which implies $X_{1}$ is uniformly distributed on $\mathcal{X}_{1}$ and $H(X_{1})=\log v$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,b)$. ∎
@@
For $F=(U_{3,4},U^{123}_{1,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
@@
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),k\neq 4,$
 $\qquad\qquad\text{ distinct }i,j,k\in N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
It holds for any $x_{4}^{\prime}\in\mathcal{X}_{4}$, which implies that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}$ , and so $H(X_{4})=\log v$ where $v=|\mathcal{X}_{4}|$.
On the other hand, the “if” part of the theorem can be proved by Lemma 2 and the fact that $a=\log{v}$ on the ray $U_{3,4}$ and whole ray $U^{123}_{1,3}$ are entropic. ∎
@@
For $F=(U_{3,4},U_{1,4})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a\geq 0,b>0$ or $(a,b)=(\log{v},0)$ for some positive integer $v$.
@@
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
@@
Consider the tripartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}$ and $(x_{i},x_{j})\in E$ if and only if $p(x_{i},x_{j})>0$, ${i,j}\in N_{3},i\neq j$. Assume $G$ has $t$ connected components. We denote the number of the vertices of $\mathcal{X}_{i}$ in the connected component $C_{j}$ by $n_{i}^{(j)},i=1,2,3,j=1,2,\cdots,t$ and the probability mass of $C_{j}$ by $p_{j}$. By Lemma 3, we obtain $n_{1}^{(j)}=n_{2}^{(j)}=n_{3}^{(j)}$ and so it can be simplified to $n^{(j)}$. The probability mass of the vertices, the edges and the triangles in $C_{j}$ are
@@
By Lemma 2 and the fact that $a=\log v$ on the ray $U_{3,4}$, and the whole ray $U_{1,4}$ are entropic, all $\mathbf{h}=(a,b)\in F$ are entropic when $a=\log v$ for positive integer $v$, and $b\geq 0$. In the following, we only need to show that for all $\mathbf{h}=(a,b)\in F$ are entropic, when $a>0,b>0$ with $a\neq\log v$ for some integer $v>0$.
Consider $a\neq\log{v},b>0$. Let $c=\lfloor 2^{a}\rfloor$. Let $n^{(j)}=c$ for all $j\in N_{t-1}$, $n^{(t)}=c+l$ for some sufficiently large $l$ and $0<p_{i}<1$, $i=1,2,\dots$, $t$ be positive values satisfying $H(p_{1},...,p_{t})=b$ such that
@@
$=(1-p_{t})\log c+p_{t}\log{(c+l)}$ (IV.E.145)
$=a.$ (IV.E.146)
Let $X^{(j)}_{1}$ $X^{(j)}_{2}$ and $X^{(j)}_{3}$ be mutually independent and uniformly distributed on $\mathbb{I}_{c}$ , and $X^{(j)}_{4}=X^{(j)}_{1}+X^{(j)}_{2}+X^{(j)}_{3}\mod c$ for $j\in N_{t-1}$. Let $X^{(t)}_{1}$, $X^{(t)}_{2}$ and $X^{(t)}_{3}$ be mutually independent and uniformly distributed on $\mathbb{I}_{c+l}$ and $X^{(t)}_{4}=X^{(t)}_{1}+X^{(t)}_{2}+X^{(t)}_{3}\mod(c+l)$. Assume $X^{(j)}_{i}$ is distributed on $\mathcal{X}^{(j)}_{i}$ for $i\in N_{4}$ and $j\in N_{t}$. Let $X_{i},i\in N_{4}$ be distributed on $\mathbb{I}_{tc+l}$ such that
$p_{X_{N_{4}}}(x_{i}+(j-1)c,i\in N_{4})$
$=$ $p_{j}p_{X^{(j)}_{1},X^{(j)}_{2},X^{(j)}_{3},X^{(j)}_{4}}(x_{1},x_{2},x_{3},x_{4})$ (IV.E.147)
for $j\in N_{t}$ and $(x_{1},x_{2},x_{3},x_{4})\in\mathcal{X}^{(j)}_{N_{4}}$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,b)$. Hence, all $\mathbf{h}=(a,b)\in F$ are entropic when $a>0,b>0$. The proof is accomplished. ∎
Theorem IV-E.5 can be considered as a corollary of [32, Lemma 4], which is restated as [36, Lemma 6].